OpenStudy (anonymous):

An acute angle θ is in a right triangle with sin θ =2/3 . What is the value of cot θ?

OpenStudy (anonymous):



OpenStudy (anonymous):

find the third side by pythagoras, you get \(\sqrt{3^2-2^2}=\sqrt{5}\) then use "adjacent over opposite"

OpenStudy (campbell_st):

you need to find the missing side in the right triangle using pythagoras' theorem  once you have found x then find \[\tan( \theta)\] lastly talke the reciprocal to find \[\cot(\theta)\]